OpenStudy (anonymous):

Evaluate the integral 3/sqrt (4-9x^2)dx, I got to the point of getting u=x du=1dx and a=sqrt(4/9) and i used the trig fcn of arcsinu/a+c but i still can't figure it out. please help!

OpenStudy (turingtest):

\[\int{3\over\sqrt{4-9x^2}}dx\]is the prob?

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

then you need a trigonometric substitution\[\frac23x=\sin\theta\]do you know how to do these from here?

OpenStudy (mimi_x3):

\[\int\limits\frac{3}{\sqrt{4-9x^{2}}} dx =>\int\limits\frac{3}{\sqrt{9(\frac{4}{9})-x^{2})}} dx\] i think that i answered this question before..

OpenStudy (turingtest):

whoops, bad sub, let me fix that...

OpenStudy (mimi_x3):

No subsitution is needed i think..

OpenStudy (turingtest):

the sub you need is\[x=\frac23\sin\theta\]@Mimi_x3 please demonstrate

OpenStudy (anonymous):

could you show me the steps? because my answer was 3arcsin(9xsqrt(4/9)/4 +c but i don't know if that is right?

OpenStudy (turingtest):

close, should be arcsin(3x/2)+C

OpenStudy (mimi_x3):

\[\frac{3}{9} \int\limits\frac{1}{\sqrt{(\frac{4}{9})-x^{2}}} dx =>\frac{1}{3} \int\limits\frac{1}{\frac{2^{2}}{3^{2}}-x^{2}} dx \] this is the step..then it's a trig inverse function..

OpenStudy (anonymous):

@mimi_x3 how did you pull out the 9 out of sq rt to get the coefficient of 3/9?

OpenStudy (turingtest):

\[\sqrt{4-9x^2}=\sqrt{9(\frac49-x^2})=3\sqrt{\frac49-x^2}\]

OpenStudy (mimi_x3):

lol, i think that something wrong..which i can't figure out at this time.

OpenStudy (turingtest):

and nice job Mimi, never would have thought to do it that way... you dropped the sqrt sign at the end though, but you still did the problem right

OpenStudy (mimi_x3):

ohh okay, small mistake my bad..